[figure: 9.7 cm]

MONTE CARLO STUDY OF PHASE DIAGRAMS OF THE MIXED SPIN-1/2 AND SPIN-1 ISING SYSTEM

D. Horváth^{a}, M. Jaščur^{a}, T. Balcerzak^{b} and A. Bobák^{a} a Department of Theoretical Physics and Geophysics, Faculty of Science, P. J. Šafárik University, Moyzesova 16, 041 054 Košice, Slovak Republic bSolid State Physics Department, University of Łódź, Pomorska 149/153, 90 236 Łódź, Poland

The standard Metropolis algorithm and histogram Monte Carlo simulations have been applied to study the phase diagrams of mixed spin-1/2 and spin-1 Ising model. Three different geometrical structures, namely, the square lattice (L × L), the bilayer system (L × L × 2) and the simple cubic lattice (L × L × L) with L ≤ 64 have been investigated and the crystal-field dependences of the critical temperature have been obtained. The Monte Carlo results are compared with those of the standard Bethe-Peierls-Weiss approximation. The occurrence of the tricritical behaviour and the first order phase transition in the system is analysed in detail and a possible relation of the model under investigation to experimental bimetallic molecular-based magnetic systems is also discussed.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : Dr. Denis Horváth

Address for correspondence : Department of Theoretical Physics and Geophysics, Faculty of Science, P. J. Šafárik University, Moyzesova 16, 041 054 Košice, Slovak Republic

Email address : horvath@kosice.upjs.sk